# kNN Classification of Malware Data Dependency Graph Features

### v1

@@ Abstract
Feature resolution impacts the ability of classifiers to make explainable inferences when applied to malware classification. We explore classification based on features constructed from data dependency graphs, and present results from k-Nearest Neighbors (kNN) classifiers. Our study demonstrates that classification based on a novel feature representation not only yields high accuracy, but also increases explainability in inference, as features of data dependency are directly representative of program behavior. We present classification results using the Microsoft Kaggle 2015 malware dataset which was processed with a novel approach to feature extraction and representation. We show that non-parametric approaches to classification in the metric space are able to obtain classification accuracy of 87.5% when applied to multi-class classification in the Kaggle malware dataset. Additionally, similarity in the metric space can be calculated directly without prior training. Our results provide evidence that data dependency graphs accurately capture both semantic and structural information.
@@ III Results
Table 2 presents accuracy metrics for the same classifier. Total accuracy is the true positive plus false positive divided by the total, in order to exclude mispredictions. In order to calculate the total accuracy we can take the true positive value of 2,314 true predictions and divide it by the size of the test set, 2,657. More detail on the total accuracy of classifiers are presented in table 3.
Figures 5 and 6 present results for kNN classifiers with k values 3 in a confusion matrix. Both classifiers have a total accuracy of 85% or better.
@@
Tables 3 and 4 present accuracy metrics for kNN classifiers with k=4 and k=5, respectively. The accuracy for each class, along with the precision, recall, specificity, and F1 score are presented.
@@ IV Discussion
One explanation for the accuracy measurements in classification is the choice of feature representation, and its ability to capture semantic information. If the data dependency graphs contained no semantic information, then we would expect the accuracy numbers to be very low. Low classification accuracy would be an indication that isomorphism of data dependency graphs between operands of instructions are not good representations of program behavior, operational semantics, or structural patterns. However, an analysis only based on operands of ’$mov$’ instructions can provide accuracy of 87.5% with non-parametric learning approaches. This indicates that the feature representation is sufficiently capturing semantic information.
@@
We highlight that our method allows for further refinements, reduction in the search space, and increases in resolution to provide more fine grained results. These approaches are recommended for future studies. In future work we intend to increase the accuracy through increasing the feature resolution, as well as explore the impact of fuzzy models including fuzzy c-means.
@@ V Conclusion
In this study we have presented results from k-Nearest Neighbors classifiers across several k values, and have evaluated these classifiers based on precision, recall, specificity, F1 score, and total accuracy. We have shown that a classifier trained on our feature representation obtained a total accuracy for a multi-class classification task of 87.5%. Similarity in the metric space can be calculated without prior training. This demonstrates the method of feature representation, and validates the hypothesis that data dependency isomorphism is representative of program behavior and operational semantics.

### v2

@@ Abstract
Explainability in classification results are dependent upon the features used for classification. Data dependency graph features representing data movement are directly correlated with operational semantics, and subject to fine grained analysis. This study obtains accurate classification from the use of features tied to structure and semantics. By training an accurate model using labeled data, this feature representation of semantics is shown to be correlated with ground truth labels. This was performed using non-parametric learning with a novel feature representation on a large scale dataset, the Kaggle 2015 Malware dataset. The features used enable fine grained analysis, increase in resolution, and explainable inferences. This allows for the body of the term frequency distribution to be further analyzed and to provide an increase in feature resolution over term frequency features. This method obtains high accuracy from analysis of a single instruction, a method that can be repeated for additional instructions to obtain further increases in accuracy. This study evaluates the hypothesis that the semantic representation and analysis of structure are able to make accurate predications and are also correlated to ground truth labels. Additionally, similarity in the metric space can be calculated directly without prior training. Our results provide evidence that data dependency graphs accurately capture both semantic and structural information for increased explainability in classification results.
@@ III Results
Figure 5 and the associated table presents accuracy metrics for the same classifier. Total accuracy is the true positive plus false positive divided by the total, in order to exclude mispredictions. In order to calculate the total accuracy we can take the true positive value of 2,314 true predictions and divide it by the size of the test set, 2,657. More detail on the total accuracy of classifiers are presented in table 3.
Figure 6 and 7 present results for kNN classifiers with k values 3 and 4 in a confusion matrix. Both classifiers have a total accuracy of 85% or better. Tables 2 and 3 present accuracy metrics for kNN classifiers with k=3 and k=4, respectively. The accuracy for each class, along with the precision, recall, specificity, and F1 score are presented.
@@ IV Discussion
One explanation for the accuracy measurements in classification is the choice of feature representation, and its ability to capture semantic information. If the data dependency graphs contained no semantic information, then we would expect the accuracy numbers to be very low. Low classification accuracy would be an indication that isomorphism of data dependency graphs between operands of instructions are not good representations of program behavior, operational semantics, or structural patterns. However, an analysis only based on operands of ’$mov$’ instructions can provide accuracy of 87.5% with non-parametric learning approaches, which can be increased by analyzing additional instructions. This indicates that the feature representation is sufficiently capturing semantic information.
@@
We highlight that our method allows for further refinements, reduction in the search space, and increases in resolution to provide more fine grained results. These approaches are recommended for future studies. In future work we intend to increase the accuracy through increasing the feature resolution, expanding the instructions analyzed, as well as explore the impact of fuzzy systems.
@@ V Conclusion
In this study we have presented results from k-Nearest Neighbors classifiers across several k values, and have evaluated these classifiers based on precision, recall, specificity, F1 score, and total accuracy. We have shown that a classifier trained on our feature representation for a single instruction obtained a total accuracy for a multi-class classification task of 87.5%, which can be repeated for additional instructions to gain an increase in accuracy. Similarity in the metric space can be calculated without prior training. This demonstrates the method of feature representation, and validates the hypothesis that data dependency isomorphism is representative of program behavior and operational semantics.